# Category-level Meta-learned NeRF Priors for Efficient Object Mapping

### v1

@@ Abstract
In 3D object mapping, category-level priors enable efficient object reconstruction and canonical pose estimation, requiring only a single prior per semantic category (e.g., chair, book, laptop). Recently, DeepSDF has predominantly been used as a category-level shape prior, but it struggles to reconstruct sharp geometry and is computationally expensive. In contrast, NeRFs capture fine details but have yet to be effectively integrated with category-level priors in a real-time multi-object mapping framework. To bridge this gap, we introduce PRENOM, a Prior-based Efficient Neural Object Mapper that integrates category-level priors with object-level NeRFs to enhance reconstruction efficiency while enabling canonical object pose estimation. PRENOM gets to know objects on a first-name basis by meta-learning on synthetic reconstruction tasks generated from open-source shape datasets. To account for object category variations, it employs a multi-objective genetic algorithm to optimize the NeRF architecture for each category, balancing reconstruction quality and training time. Additionally, prior-based probabilistic ray sampling directs sampling toward expected object regions, accelerating convergence and improving reconstruction quality under constrained resources. Experimental results on a low-end GPU highlight the ability of PRENOM to achieve high-quality reconstructions while maintaining computational feasibility. Specifically, comparisons with prior-free NeRF-based approaches on a synthetic dataset show a 21% lower Chamfer distance, demonstrating better reconstruction quality. Furthermore, evaluations against other approaches using shape priors on a noisy real-world dataset indicate a 13% improvement averaged across all reconstruction metrics, a boost in rotation estimation accuracy, and comparable translation and size estimation performance, while being trained for 5× less time.
@@ IV-C Comparison with category-level shape prior approaches
Metrics To evaluate reconstruction quality, we use CD and CR as before. Moreover, we also incorporate Earth Mover’s Distance (EMD) to provide a more robust assessment of global shape structure, since NeRF-based representations tend to produce sharper local features than DeepSDF-based representations, and CD alone can be biased toward these details. For object state estimation, we compute translation, rotation, and size errors.
Results and discussion PRENOM achieves an improvement in reconstruction quality of approximately 13% averaged across all metrics compared to the second-best approach, as shown in Table II. Notably, it outperforms other methods in reconstruction quality while only being trained for 5$\times$ less time and with a 30% smaller prior size, underscoring its efficiency. Additionally, comparisons with our prior-free version re-emphasize that our prior-based approach not only enables canonical object state estimation but also significantly enhance reconstruction quality of NeRF-based object representations.
@@
Moreover, for object state estimation, our prior-density-based pose canonicalization significantly improves rotation accuracy (as shown in Table III), as ¯$\overline{\mathcal{F}}_{\mathbf{CHAIR}}$ allows reasonable scoring for chairs of varying geometry. However, unlike other methods that jointly optimize object state and shape reconstruction, PRENOM only estimates the state once and thus exhibits higher translation and size errors. Notably, our size estimates are directly derived from the noisy point cloud extents. In NeRF-based representations, the estimated size only needs to sufficiently bound the expected mesh region, often leading to overcompensation, as is the case here.
@@ V Conclusion
Despite its advantages, PRENOM is inherently limited by its NeRF-based representation, which struggles with out-of-view feature reconstruction. While our priors help mitigate these issues, the absence of hard constraints to enforce consistency eventually leads to noisier reconstructions, especially if trained for extended periods. Additionally, object pose is not refined during training, resulting in slightly worse pose estimation than some existing shape prior methods. Future work will focus in these directions.

### v2

@@ Abstract
In 3D object mapping, category-level priors enable efficient object reconstruction and canonical pose estimation, requiring only a single prior per semantic category (e.g., chair, book, laptop). Recently, DeepSDF has predominantly been used as a category-level shape prior, but it struggles to reconstruct sharp geometry and is computationally expensive. In contrast, NeRFs capture fine details but have yet to be effectively integrated with category-level priors in a real-time multi-object mapping framework. To bridge this gap, we introduce PRENOM, a Prior-based Efficient Neural Object Mapper that integrates category-level priors with object-level NeRFs to enhance reconstruction efficiency while enabling canonical object pose estimation. PRENOM gets to know objects on a first-name basis by meta-learning on synthetic reconstruction tasks generated from open-source shape datasets. To account for object category variations, it employs a multi-objective genetic algorithm to optimize the NeRF architecture for each category, balancing reconstruction quality and training time. Additionally, prior-based probabilistic ray sampling directs sampling toward expected object regions, accelerating convergence and improving reconstruction quality under constrained resources. Experimental results on a low-end GPU highlight the ability of PRENOM to achieve high-quality reconstructions while maintaining computational feasibility. Specifically, comparisons with prior-free NeRF-based approaches on a synthetic dataset show a 21% lower Chamfer distance, demonstrating better reconstruction quality. Furthermore, evaluations against other approaches using shape priors on a noisy real-world dataset indicate a 13% improvement averaged across all reconstruction metrics, and comparable pose and size estimation accuracy, while being trained for 5× less time.
@@ IV-C Comparison with category-level shape prior approaches
Metrics To evaluate reconstruction quality, we use CD and CR as before. Moreover, we also incorporate Earth Mover’s Distance (EMD) to provide a robust assessment of global shape, since NeRF-based approaches tend to produce sharper local features than DeepSDF-based approaches, and CD alone can be biased toward these details.
For object state estimation, we compute translation, rotation, and size errors. Morover, similar to PRENOM, the baseline methods also assume that chairs lie parallel to the ground to compute the initial pose. However, since these methods update the pose during reconstruction, we align their final poses’ vertical axes with the ground truth before evaluation to ensure fairness.
Results and discussion PRENOM achieves an improvement in reconstruction quality of approximately 13% averaged across all metrics compared to the second-best approach, as shown in Table II. Notably, it outperforms other methods in reconstruction quality while only being trained for 5$\times$ less time and with a 30% smaller prior size, underscoring its efficiency. Additionally, comparisons with our prior-free version re-emphasize that our prior-based approach not only enables canonical object state estimation but also significantly enhance reconstruction quality.
@@
Moreover, despite not jointly optimizing object state with reconstruction, our method achieves comparable object state recovery results as can be seen in Table III. We do not consider the improvement in rotation accuracy to be significant, as PRENOM does not adjust pose during reconstruction and thus benefits from the initial assumption that objects are parallel to the ground. While this initial assumption is applied to all methods, the baseline approaches are more susceptible to noise as they optimize pose during the reconstruction process, which negatively impacts the rotation accuracy.
@@ V Conclusion
Despite its advantages, PRENOM is inherently limited by its NeRF-based representation, which struggles with out-of-view feature reconstruction. Although our priors help mitigate these issues, the absence of hard constraints to enforce consistency eventually leads to noisier reconstructions, especially if trained for extended periods. Additionally, object state is not refined during training, resulting in slightly worse state estimation than some existing shape prior methods. Future work will focus in these directions.